solve for x x=
9(5x-3) +5(2x+4)=2(3x+3)

Answer by AnlytcPhil(1806)

solve for x  

9(5x - 3) + 5(2x + 4) = 2(3x + 3)

--------------------------------------

Multiply the 9 by the 5x, getting 45x.  So far we have:

45x

--------------------------------------

9(5x - 3) + 5(2x + 4) = 2(3x + 3)

Multiply the 9 also by the -3.  So far we have:

45x - 27

--------------------------------------

9(5x - 3) + 5(2x + 4) = 2(3x + 3)

Multiply the +5 by the 2x.  So far we have:

45x - 27 + 10x

--------------------------------------

9(5x - 3) + 5(2x + 4) = 2(3x + 3)

Multiply the +5 also by the +4.  So far we have:

45x - 27 + 10x + 20

--------------------------------------

9(5x - 3) + 5(2x + 4) = 2(3x + 3)

Copy the equal sign:

So far we have:

45x - 27 + 10x + 20 =

--------------------------------------

9(5x - 3) + 5(2x + 4) = 2(3x + 3)

Multiply the 2 by the 3x.  So far we have:

45x - 27 + 10x + 20 = 6x

--------------------------------------

9(5x - 3) + 5(2x + 4) = 2(3x + 3)

Multiply the 2 also by the +3.  So far we have:

45x - 27 + 10x + 20 = 6x + 6

--------------------------------------

45x - 27 + 10x + 20 = 6x + 6

Now combine the 45x and the +10x. So far we have

      55x - 27 + 20 = 6x + 6

Now combine the -27 and the +20:

            55x - 7 = 6x + 6

Now get rid of the -7 on the left by adding 
its opposite +7 to both sides of the equation:

            55x - 7 = 6x + 6
                + 7      + 7
          --------------------
            55x     = 6x + 13

Close up the gap:

                55x = 6x + 13

Now get rid of the 6x on the right by adding 
its opposite -6x to both sides of the equation:

                55x = 6x + 13
               - 6x  -6x 
             ------------------
                49x =      13

Close up the gap

                49x = 13

Get rid of the 49 coefficient by dividing both sides
by 49

                49x     13
               ----- = ----
                 49     49

Cancel the 49's on the left

                 1  
                49x     13
               ----- = ----
                 49     49
                  1

                        13
                   x = ----
                        49

That's the answer: x = 13/49, thirteen forty-ninths.

Edwin